OpenStudy (anonymous):

A solid cylinder cut into two equal parts and placed side by side. If the diameter of the cylinder is 7cm and the height is 8cm, find the total surface area of the two parts. (Answer given in the book is 365cm square)

OpenStudy (anonymous):

do you know how to find surface area of the full cylinder

OpenStudy (anonymous):

\[full _{SA}+rectangles_{SA}=2 \pi r^2+2\pi rh=2\pi \frac{7}{2}(\frac{7}{2}+8)+2(7)(8)\]

OpenStudy (anonymous):

I know the formula is 2pi r(r+h)

OpenStudy (anonymous):

yes so how much you get when you put radius r=7/2 and h=8

OpenStudy (anonymous):

Sorry I duno how you get this figure from the formula u stated 2pirsquare + 2pirh = 2pi7/2(7/2+8) + 2(7)(8) but what i get is 2pi7/2square + 2pi7/2(8) = 253 Can you explain this please? Thanks so much for your time

OpenStudy (anonymous):

on the answer you gt you are supposed to add 2(7)(8)

OpenStudy (anonymous):

253+112=365

OpenStudy (anonymous):

oops ok i understand already thanks! ^^